OpenStudy (anonymous):

how do I simplify 2/ square root of 5?

OpenStudy (anonymous):

This is called "rationalizing" a denominator (he square root of any non perfect square is irrational). You can do this by moving it to the top. If you don't know how: What is a/a? Well, anything (save 0) divided by itself is 1, so it's 1. What's x * 1? That's easy: it's x because anything multiplied by itself is one. Combine those two. 2/sqrt(5) * b/b =2/sqrt(5) for all non zero b. What if we make b = sqrt(5)? What happens? 2/sqrt(5) * sqrt(5)/sqrt(5) = 2sqrt(5)/5 because a squareroot times itself (aka squared) cancels out. Answer is 2sqrt(5)/5.

OpenStudy (campbell_st):

\[2/\sqrt{5} \times \sqrt{5}/\sqrt{5}\]